How would you put y+1=0 into slope intercept form? ( y=mxtb)
Answer by algebrahouse.com(1659)

How would you put y+1=0 into slope intercept form?

Slope-intercept form is y = mx + b
m is the slope
b is the y-intercept

y + 1 = 0
y = -1 {subtracted 1 from both sides....this is slope-intercept form}

this would be a horizontal line passing through -1 on the y-axis
the slope of this line is 0

For more help from me, visit: www.algebrahouse.com